OpenStudy (anonymous):

21-2x=5y My love for you is like a concave up function because it is always increasing.

OpenStudy (texaschic101):

what is your question ?

OpenStudy (anonymous):

solve for y and x

OpenStudy (texaschic101):

21 - 2x = 5y --- divide both sides by 5 21/5 - 2/5x = y 21 - 2x = 5y --- subtract 21 from both sides -2x = 5y - 21 --- divide both sides by -2 x = -5/2y + 21/2

OpenStudy (anonymous):

yup I was right

OpenStudy (texaschic101):

good to hear...confirmation is always good :)

OpenStudy (anonymous):

mhm :)